string of violin 36 cm long and mass of.2 gram at what tension it must be stretched to tune 100Hz

1
Expert's answer
2021-01-10T18:10:34-0500

"T=4f^2ml/n^2=4\cdot 100^2\cdot 0.002\cdot 0.36/1^2=28.8(N)"